A grocer wishes to mix $1.20 tea with $1.50 tea to make a mixture of 60 pounds worth $1.30 a pound. How many pounds of each kind must he mix?
Answer by richwmiller(17219)

You had $120 tea. I assume you mean $1.20
let x= tea at 1.20
x*120+(60-x)*150=60*130
x-40
60-x=20